# Dynamics of Spinning Particles in Reissner-Nordström black hole exterior

### v1

@@ I Introduction
Motivated by EMRIs, the closed-form solutions for the motion of a spinning particle near a spherically symmetric black hole are studied in Witzany and Piovano (2024) to leading order in a small spin. These analytical solutions may have applications to the generated gravitational waveforms arising from EMRIs LISA Consortium Waveform Working Group (2023); Drummond and Hughes (2022a, b) as well as the understanding of black hole accretion Fabian et al. (2020); Page and Thorne (1974); Reynolds and Begelman (1997); Schnittman et al. (2016); Jai-Akson et al. (2017). The motion of a spinless particle around Kerr family black holes along geodesics has been extensively studied in the literature Gralla and Lupsasca (2020) and by us Wang et al. (2022); Li et al. (2023); Ko et al. (2024). The solutions of the trajectories can be written in terms of elliptical integrals and Jacobi elliptic functions Abramowitz and Stegun (1964), in which the orbits are manifestly real functions of the Mino time Mino (2003) and also the initial conditions can be explicitly specified Gralla and Lupsasca (2020). Our current work will be extended to a spinning particle but around the exterior of a charged black hole with the Reissner-Nordström (RN) metric by following Witzany and Piovano (2024). For a small spin, we will first derive the radial potential from the underlying equations of motion and classify its roots to construct the parameter space diagram for various types of orbits. The radius of the innermost stable circular orbits (ISCOs) with the spin effect will be explored Jefremov et al. (2015); Zhang et al. (2018). Then the closed-form solutions for the motion under consideration will be obtained for both bound and unbound orbits of a particle with an aligned and misaligned spins with respect to its orbital angular momentum. Theoretical considerations, together with recent observations of structures near Sagittarius A* by the GRAVITY experiment Abuter et al. (2018) (GRAVITY Collaboration), indicate the possible presence of a small electric charge of the central supermassive black hole Zajaček et al. (2018, 2019). Thus, it is of great interest to explore the particle orbiting around a charged black hole. One of the motions to be studied is an orbit traveling between two turning points along the radial direction, which is a generalization of Witzany and Piovano (2024) to a charged black hole. Another interesting orbit is that the particle starts from near ISCO and inspirals into the black hole Mummery and Balbus (2022, 2023a); Ko et al. (2024). Lastly, we will also consider the homoclinic orbit, which is the separatrix between bound and plunging orbits and a solution that asymptotes to an energetically bound, unstable spherical orbit Levin and Perez-Giz (2009); Li et al. (2023). The existence of the homiclinic orbit is relevant to the chaotic behavior Suzuki and Maeda (1997) although there is some evidence that terms linear in spin do not cause the emergence of chaos, which in turn has little effect on the overall dynamics of an EMRI with a non-spinning black hole Zelenka et al. (2020). The unbound motion starting from space infinity, traveling towards a black hole, hitting the turning point, and returning to space infinity is also shown in Appendix B for comparison.
@@ II Equations of motion for a spinning particle
where $f\left(r\right)=\frac{1}{h\left(r\right)}=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}$ with black hole mass $M$ and charge $Q$. For the case of $Q=0$, the metric is reduced to that of the Schwarzschild black hole. The Killing vectors of this metric are generators of time translation and rotations around the $x,y,z$ axes
@@
with the constant of motion in the form Drummond and Hughes (2022a)
@@
This is called the orbital angular momentum 4-vector Drummond and Hughes (2022a). It can be shown that (30) is parallel transported along the geodesics in general spherically symmetric static spacetimes due to (29). As a result, for the spinning particle following the non-geodesics through the curvature and spin interaction, $\frac{Dl^{\mu}}{d\sigma_{m}}\propto\mathcal{O}(s)$. The aligned component of spin with $l^{\mu}$ is an approximate constant of motion to order in linear spin
@@
To expand the spin $s^{\mu}$ in (7) in terms of the components, we introduce the tetrad $\{e_{(0)}^{\mu},e_{(1)}^{\mu},e_{(2)}^{\mu},e_{(3)}^{\mu}\}$ from Marck (1983), which are orthogonal to each other and also parallel transported along the geodesics. We start from the first leg $e_{(0)}^{\mu}$ to be the 4-velocity of the geodesic with $\dot{\vartheta}=0$ and $\vartheta=\pi/2$, given from (34), (35), and (36) with $s_{\lVert}=0$, as
@@
$\tilde{e}^{\mu}_{(2)}$ $=\left(\frac{\gamma_{m}\lambda_{m}}{f\sqrt{\lambda_{m}^{2}+r^{2}}},\frac{\lambda_{m}\dot{r}}{r},0,\frac{\sqrt{\lambda_{m}^{2}+r^{2}}}{r^{2}}\right)\,$ (44)
@@
$s^{r}$ $=\sqrt{s^{2}-s_{\parallel}^{2}}\left(\frac{\gamma_{m}r\sin\psi}{\sqrt{fh(\lambda_{m}^{2}+r^{2})}}+\frac{\lambda_{m}\dot{r}\cos\psi}{r}\right),$ (49)
@@ III Roots of the radial potential and innermost stable circular orbit (ISCO)
Figure 1: Diagram of parameter space $(\lambda_{m},\gamma_{m})$ with fixed charge $Q/M=0.5$ and small spin $s_{\parallel}/m=0.1$ for (a) the bound motion $\gamma_{m}<1$ in Sec. IV and (b) the unbound motion $\gamma_{m}>1$ in Appendix B. The blue (red) line shows the parameters for the stable (unstable) double root of $r_{m3}=r_{m4}$ for bound (unbound) motion and the red line for the unstable double root of $r_{m2}=r_{m3}$ for bound motion in Sec. IV.2. For bound motion, two double roots merge at the triple root at the point D in Sec. IV.3. The light yellow region is for the parameters with four distinct real-valued roots in Sec. IV.1. The light green (red) region is for the parameters with a pair complex-conjugate roots, $r_{m3}=r_{m4}^{*}$ ($r_{m3}=r_{m2}^{*}$) and two real-valued roots $r_{m2}>r_{m1}$ ($r_{m4}>r_{m1}$).
@@ IV.1 Oscillatory orbit between two turning points
${\Upsilon^{B}_{\tau_{m}}=am\left(X^{B}(\tau_{m})\mid k^{B}\right)=\nu_{r_{i}}\sin^{-1}\left(\sqrt{\frac{(r_{m4}-r_{m2})(r(\tau_{m})-r_{m3})}{(r_{m4}-r_{m3})(r(\tau_{m})-r_{m2})}}\right),}$ (99)
@@
in (62) with the obtained functions above. The periods of oscillation along the coordinates $r$ and $\delta\vartheta$ are some of the main results of this work. They will provide useful information for studying motion in the frequency domain Drummond and Hughes (2022b); Witzany and Piovano (2024). According to Drummond and Hughes (2022a, b); Witzany and Piovano (2024), the most straightforward analytical solutions together with the period of oscillation of the orbits may find their application to couple them to a Teukolsky solver to generate fluxes of gravitational waves sourced by the motion in a way to facilitate the Teukolsky equation computation.

### v2

@@ I Introduction
Motivated by EMRIs, the closed-form solutions for the motion of a spinning particle near a spherically symmetric black hole are studied in Witzany and Piovano (2024) to leading order in a small spin. These analytical solutions may have applications to the generated gravitational waveforms arising from EMRIs LISA Consortium Waveform Working Group (2023); Drummond and Hughes (2022a, b) as well as the understanding of black hole accretion Fabian et al. (2020); Page and Thorne (1974); Reynolds and Begelman (1997); Schnittman et al. (2016); Jai-Akson et al. (2017). The motion of a spinless particle around Kerr family black holes along geodesics has been extensively studied in the literature Gralla and Lupsasca (2020); Fujita and Hikida (2009) and by us Wang et al. (2022); Li et al. (2023); Ko et al. (2024). The solutions of the trajectories can be written in terms of elliptical integrals and Jacobi elliptic functions Abramowitz and Stegun (1964), in which the orbits are manifestly real functions of the Mino time Mino (2003) and also the initial conditions can be explicitly specified Gralla and Lupsasca (2020). In Kerr spacetime, and even for the trajectories of spinning particles, there exist a number of constants of motion in the linear spin order, that makes this particle’s dynamics integrable Rüdiger (1981, 1983). Moreover, the separation of the Hamilton-Jacobi equation for the motion of spinning test particles and the corresponding analytical solution have been studied in Witzany (2019); Skoupý and Witzany (2024). Our current work considers a spinning particle but orbiting around the exterior of a charged black hole with the Reissner-Nordström (RN) metric by following Witzany and Piovano (2024) to explicitly work out the analytical solutions. For a small spin, we will first derive the radial potential from the underlying equations of motion and classify its roots to construct the parameter space diagram for various types of orbits. The radius of the innermost stable circular orbits (ISCOs) with the spin effect will be explored Jefremov et al. (2015); Zhang et al. (2018). Then the closed-form solutions for the motion under consideration will be obtained for both bound and unbound orbits of a particle with an aligned and misaligned spins with respect to its orbital angular momentum. Theoretical considerations, together with recent observations of structures near Sagittarius A* by the GRAVITY experiment Abuter et al. (2018) (GRAVITY Collaboration), indicate the possible presence of a small electric charge of the central supermassive black hole Zajaček et al. (2018, 2019). Thus, it is of great interest to explore the particle orbiting around a charged black hole. One of the motions to be studied is an orbit traveling between two turning points along the radial direction, which is a generalization of Witzany and Piovano (2024) to a charged black hole. Another interesting orbit is that the particle starts from near ISCO and inspirals into the black hole Mummery and Balbus (2022, 2023a); Ko et al. (2024); Dyson and Van De Meent (2023). Lastly, we will also consider the homoclinic orbit, which is the separatrix between bound and plunging orbits and a solution that asymptotes to an energetically bound, unstable spherical orbit Levin and Perez-Giz (2009); Li et al. (2023); Stein and Warburton (2020); Ng and Teo (2025). The existence of the homiclinic orbit is relevant to the chaotic behavior Suzuki and Maeda (1997) although there is some evidence that terms linear in spin do not cause the emergence of chaos, which in turn has little effect on the overall dynamics of an EMRI with a non-spinning black hole Zelenka et al. (2020). The unbound motion starting from space infinity, traveling towards a black hole, hitting the turning point, and returning to space infinity is also shown in Appendix B for comparison. The Mathematica codes for all calculations are provided by us in Ciou et al. (2025).
@@ II Equations of motion for a spinning particle
where $f\left(r\right)={1}/{h\left(r\right)}=1-{2M}/{r}+{Q^{2}}/{r^{2}}$ with black hole mass $M$ and charge $Q$. For the case of $Q=0$, the metric is reduced to that of the Schwarzschild black hole. The Killing vectors of this metric are generators of time translation and rotations around the $x,y,z$ axes
@@
with the constant of motion in the form Dixon (1970)
@@
This is called the orbital angular momentum 4-vector Drummond and Hughes (2022a). It can be shown that (30) is parallel transported along the geodesics in general spherically symmetric static spacetimes due to (29). As a result, for the spinning particle following the non-geodesics through the curvature and spin interaction, ${Dl^{\mu}}/{d\sigma_{m}}\propto\mathcal{O}(s)$. The aligned component of spin with $l^{\mu}$ is an approximate constant of motion to order in linear spin
@@
To expand the spin $s^{\mu}$ in (7) in terms of the components, we introduce the tetrad $\{e_{(0)}^{\mu},e_{(1)}^{\mu},e_{(2)}^{\mu},e_{(3)}^{\mu}\}$ from Witzany and Piovano (2024), which are orthogonal to each other and also parallel transported along the geodesics. We start from the first leg $e_{(0)}^{\mu}$ to be the 4-velocity of the geodesic with $\dot{\vartheta}=0$ and $\vartheta=\pi/2$, given from (34), (35), and (36) with $s_{\lVert}=0$, as
@@
$\tilde{e}^{\mu}_{(2)}$ $=\left(\frac{\gamma_{m}\lambda_{m}}{f\sqrt{\lambda_{m}^{2}+r^{2}}},\frac{\lambda_{m}\dot{r}}{\sqrt{\lambda_{m}^{2}+r^{2}}},0,\frac{\sqrt{\lambda_{m}^{2}+r^{2}}}{r^{2}}\right)\,$ (44)
@@
$s^{r}$ $=\sqrt{s^{2}-s_{\parallel}^{2}}\left(\frac{\gamma_{m}r\sin\psi}{\sqrt{fh(\lambda_{m}^{2}+r^{2})}}+\frac{\lambda_{m}\dot{r}\cos\psi}{\sqrt{\lambda_{m}^{2}+r^{2}}}\right),$ (49)
@@ III Roots of the radial potential and innermost stable circular orbit (ISCO)
Figure 1: Diagram of parameter space $(\lambda_{m},\gamma_{m})$ with fixed charge $Q/M=0.5$ and small spin $s_{\parallel}/m=0.1$ for (a) the bound motion $\gamma_{m}<1$ in Sec. IV and (b) the unbound motion $\gamma_{m}>1$ in Appendix B. The blue line shows the parameters for the stable (unstable) double root of $r_{m3}=r_{m4}$ for bound (unbound) motion and the red line for the unstable double root of $r_{m2}=r_{m3}$ for bound motion in Sec. IV.2. For bound motion, two double roots merge at the triple root at the point D in Sec. IV.3. The light yellow region is for the parameters with four distinct real-valued roots in Sec. IV.1. The light green (red) region is for the parameters with a pair complex-conjugate roots, $r_{m3}=r_{m4}^{*}$ ($r_{m3}=r_{m2}^{*}$) and two real-valued roots $r_{m2}>r_{m1}$ ($r_{m4}>r_{m1}$).
@@ IV.1 Oscillatory orbit between two turning points
${\Upsilon^{B}_{\tau_{m}}={\rm am}\left(X^{B}(\tau_{m})\mid k^{B}\right)=\nu_{r_{i}}\sin^{-1}\left(\sqrt{\frac{(r_{m4}-r_{m2})(r(\tau_{m})-r_{m3})}{(r_{m4}-r_{m3})(r(\tau_{m})-r_{m2})}}\right),}$ (99)
@@
in (62) with the obtained functions above. The periods of oscillation along the coordinates $r$ and $\delta\vartheta$ are some of the main results of this work. They will provide useful information for studying motion in the frequency domain Drummond and Hughes (2022b); Witzany and Piovano (2024). According to Drummond and Hughes (2022a, b); Witzany and Piovano (2024), the analytical solutions, together with the obtained period of oscillation of the orbits, may find their application to couple them to a Teukolsky solver to generate gravitational waves sourced by the motion, thus facilitating the Teukolsky equation computation LISA Consortium Waveform Working Group (2023).